OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (8, 12) and has a slope of −2? y = x +

OpenStudy (anonymous):

plug in -2 for m and place the 12 in place of y and 8 for x and solve for b: 12=-2(8)+b 12=-16+b (add 16 to both sides) 28=b. now rewrite the equation